A mountain lion decides to leap across a 5.00 m wi
Answer #1

6) time taken to reach opposite edge = (5+2)/(10 cos 30 degree)

= 0.8082 s

vertical height =ut-0.5gt^2 =  5*0.8082 -4.9*0.8082^2

= 0.8 m

vertical height on opposide side = 0.8+4 = 4.8 m

option b is correct

7) using second equation of motion

- 4 = 5t - 4.9t^2

t = 1.548 s

excess distance =  10*cos 30 degree *1.548 - 7

= 6.4 m

option e is correct